an ammetor is reads upto 1ampere its internal resistance is 0.81 ohm .to increase the range to 10 ampere  the value of required shunt is
 

Grade:12th pass

1 Answers

Khimraj
5 years ago
Resistance of ammeter Rg = 0.81ohm
ammeter current Ig = 1A.
New ammeter range = 10A.
shunt resistance Rs =?
Rs = (Ig/(I-Ig))Rg = (1/9)*0.81 = 0.09 ohm.
Hope it clears. If you like anser please approve it.